OpenStudy (anonymous):

For each function, what is the output of the given input? For f(x) = 4x + 7, find f(4)

OpenStudy (debbieg):

Just plug in 4 where you have an x in the equation, and simplify.

OpenStudy (debbieg):

f(x) = 4x + 7 so: f(4)=4*4 + 7 = ?

OpenStudy (anonymous):

f(x) = 4*4+7 f(x) = 16+ 7=23

OpenStudy (debbieg):

right :)

OpenStudy (anonymous):

So that is all there is to it?

OpenStudy (debbieg):

yup :) easy peasy, huh? :)

OpenStudy (anonymous):

Yes thank you :)